The equation x2−8 x+y2− 14 y+z2+ 16 z= -113 represents a sphere with radius Blank 1. Calculate the answer by read surrounding text.



and center ( Blank 2. Calculate the answer by read surrounding text.



, Blank 3. Calculate the answer by read surrounding text.



, Blank 4. Calculate the answer by read surrounding text.



).





1
Expert's answer
2022-03-14T04:49:34-0400

x2 − 8x + y2 − 14y + z2 + 16z = -113

Let reform this equation in a form:

(x - a)2 + (y - b)2 + (z - c)2 = R2, (1)

where R would be radius of the sphere and (a, b, c) - its center.

(x2 − 8x + 16) + (y2 − 14y + 49) + (z2 + 16 z + 64) = -113 + 16 + 49 + 64;

(x - 4)2 + (y - 7)2 + (z + 8)2 = 16;

(x - 4)2 + (y - 7)2 + (z + 8)2 = 42.

From the formula (1) we can say, that this equation represents a sphere with radius R = 4 and center (4, 7, -8).